OpenStudy (anonymous):

*Use a quadratic equation to solve : 1. the larger of two numbers is 5 more than the smaller. The product of the numbers is 204. Find the numbers. 2 . find two positive numbers in the ratio 2:3 whose product is 2400.

OpenStudy (pfenn1):

If x is the larger number and y is the smaller number, what would be the equation for "the larger of two numbers is 5 more than the smaller"?

OpenStudy (pfenn1):

How about "the product of the numbers is 204"?

OpenStudy (anonymous):

idk that's why i am asking, i am confused with this!

OpenStudy (pfenn1):

If x (the larger of two numbers) is 5 more than y (the smaller), how would write that in an equation?

OpenStudy (anonymous):

x>5y ?

OpenStudy (pfenn1):

Close, what you wrote could be written as the larger number is greater than 5 times the smaller number. x (the larger of two numbers) is 5 more than y (the smaller) x = 5 + y

OpenStudy (anonymous):

so it's x=5+y ?

OpenStudy (pfenn1):

Yes. A lot of what you are trying to do in "word problems" is to try to develop the equations that underlie the sentences. Do you know which mathematical function is implied by the "product"?

OpenStudy (anonymous):

multiplyinng?

OpenStudy (pfenn1):

Correct. so what would "the product of the numbers is 204" imply?

OpenStudy (anonymous):

204=x? (5+y) * y = 204?

OpenStudy (pfenn1):

Yes. The first equality 204=x? should be 204=xy but the second is correct. (5+y) * y = 204 So now rearrange and solve for y. And then use x=5+y to solve for x.

OpenStudy (anonymous):

what about the second one can you help me with it?

OpenStudy (anonymous):

y= -17 or 12

OpenStudy (pfenn1):

Sure. If x and y are two positive numbers, then "two numbers in the ratio 2:3" I think implies that \[\left( x \over y \right) = \left( 2 \over 3 \right)\] The second equation is based on the "product" of the two numbers is 2400. Can you write that second equation?

OpenStudy (pfenn1):

Yes, I got the same solution to question 1. y=-17 or 12

OpenStudy (anonymous):

\[\left(\begin{matrix}2 \\ 3\end{matrix}\right) = \left(\begin{matrix}2400 \\ x\end{matrix}\right)\] which means x= 3600?

OpenStudy (pfenn1):

I'm not sure what you did there... The "product" of the two numbers is 2400 implies \[xy=2400\]If you rearrange the first equation, \[\left( x \over y \right) = \left( 2 \over 3 \right)\]To solve for x you get\[x = \left( 2y \over 3 \right)\]Plug that into \[xy=2400\]and solve for y

OpenStudy (anonymous):

so it is 3600

OpenStudy (pfenn1):

No, that isn't what I got. \[\left( 2y ^{2}\over 3 \right)=2400\]rearrange to get\[y ^{2}=2400 \left( 3 \over 2 \right)=3600\]

OpenStudy (pfenn1):

3600 can't be the right answer because 3600x=2400 would imply that x is <<<3600 and the ratio of x/y = 2/3. Sometimes you can check your answers by just seeing if they make any sense.

OpenStudy (pigrenex):

try to solve it by these two equations \[x \div y=2\div3\] and \[x \times y =2400\] rearrange the equation 2, isolating y. finally, insert it in the equation 1